evaluate ∫ (sec^2 (x) + cosec^2 (x)) dx

evaluate
∫ (sec^2 (x) + cosec^2 (x)) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

I = \int (sec^{2}x+cosec^{2}x)dx
I = \int sec^{2}xdx+\int cosec^{2}xdx
I = tanx+\int cosec^{2}xdx
I = tanx+(-cotx) + c
I = tanx - cotx + c